OpenStudy (anonymous):

a rectangle has a width of 2 inches and contains three inscribed circles calculate the area inside the rectangle that is not inside any of the circles

OpenStudy (anonymous):

what a question?

OpenStudy (anonymous):

find erea of rectangle without 3 circle?

OpenStudy (anonymous):

yes please help

OpenStudy (anonymous):

do you need a picture of it?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Is it like [OOO] ?

OpenStudy (anonymous):

exactly

OpenStudy (anonymous):

r=2/2=1 A=1^2pi 3circle = 3pi rectangle leng = 6 A recta =2*6=12 A=12-3pi A=2.575222039 u round

OpenStudy (anonymous):

Okay, the width of the rectangle is 2. We need to figure out the height. Since there are 3 circles which fit exactly into this width of 2, the diameter of each circle must be 2/3 inches By this we know that the height of the rectangle is 2/3 inches. So the area of the rectangle is 2 x (2/3) = 4/3 inches squared. subtract from that area, the area of each circle. area of the circle is pi (2/6)^2

OpenStudy (anonymous):

Oh i guess it depends if they mean 2 inches is the short or long length of the rectangle heh, I assumed long